How do you solve 4u-6 greater than or equal to 6u-20
Answer by 2897696(96)

>= means greater or equal to <= means less than or equal to
4u-6 >= 6u-20
+6 >= +6
------------
4u>=6u-14
-6u >=-6u
------------
-2u>=-14
switch sign (flip sign because multiplied negative number to the other side)
-2u<=-14
---<=---
-2 <= -2
u=7
answer: u=7